A study of corruption for a certain geographic region showed an average of 5 corruptions occur per 20,000 people. In a city of 80,000 people, find the probability that at least 3 corruptions occur.
1
Expert's answer
2020-10-26T19:08:44-0400

Poisson distribution

"\lambda=5×4=20"

"Px\ge3)=1-P(x\le2)"

"=1-(\frac{\lambda^xe^{-\lambda}}{x!}),x=0,1,2"

"=1-(\frac{20^0e^{-20}}{0!}+\frac{20^1e^{-20}}{1!}+\frac{20^2e^{-20}}{2!})"

"=1-(221e^{-20})"

"=0.9999999979"